- My blog
- Post new blog entry
- Top bloggers
- Recent posts
Fiber bundles pulled away from magnet busbars
Updated on Sat, 2018-03-10 16:32. Originally created by adams92 on 2018-03-10 16:32.We needed had previously used zip ties with small labels as stand offs to keep the fiber bundles away from the magnet busbars, since they get hot (~130 F when measured with an IR thermometer) (s
Mapping Tests March 9, 2018
Updated on Sat, 2018-03-10 16:15. Originally created by adams92 on 2018-03-10 16:15.Attached are the seven pattern file outputs confirming our mapping just before the run.
Vped Scan 2018 Take 2
Updated on Sat, 2018-03-10 00:45. Originally created by rjreed on 2018-03-10 00:29.Since some boards had been swapped since the last vped scan, the procedure was repeated.
The runs are:
19068022 CAMAC OFF EPD_TUFFmap.txt
NICA Days 2017 proceedings v2
Updated on Fri, 2018-03-09 19:12. Originally created by siejka on 2018-03-09 19:12.NICA Days 2017 proceedings v2
Production priority status
Updated on Fri, 2018-03-30 06:37 by testadmin. Originally created by jeromel on 2018-03-09 15:29.Overview of recent heavy flavor results at STAR
Title: Overview of recent heavy flavor results at STAR
Conference: Excited QCD
Time: 20+10 min
EPD First beam data of 2018!
Updated on Fri, 2018-03-09 08:21. Originally created by rjreed on 2018-03-09 08:19.I have to run to teach class, but wanted to throw up some plots:
Run 19067069 ADC spectra:
Run17 J/Psi Search
Updated on Thu, 2018-03-08 11:12. Originally created by heppel on 2018-03-08 11:12.
Radiation Damage Simulation Double Check
Updated on Thu, 2018-03-08 00:35. Originally created by bsumma on 2018-03-08 00:35.Checking for possible causes for radiation damage simulation for small cells not matching data.
Cosmic prepost epd 2018 Part III
Updated on Wed, 2018-03-07 16:57. Originally created by rjreed on 2018-03-07 16:38.The goal is to make sure we are in the correct crossing.
Analyzing bbc-tof triggered events from 19063026 - 19063058, we saw the following distribution